### v2

# A mechanical characterization of CMC surfaces

Abstract.
The speed of a ball rolling without skidding or spinning on a surface $S$ is the length of the velocity of its center, or, equivalently, the length of its angular velocity vector. We show that if the speed only depends on $p\in S$, then $S$ has nonzero constant mean curvature; and, conversely, that if the mean curvature of $S$ is constant and equal to $H\neq 0$, then either $S$ is a sphere or the ball of radius $1/H$ rolls on $S$ with direction-independent speed. It follows that the only surfaces where the speed is constant are subsets of planes, circular cylinders, and spheres.
Key words and phrases:
Constant mean curvature, constant principal curvatures, Darboux frame, extrinsic rolling, kinematics of rolling motions
2020 Mathematics Subject Classification:
Primary 53A05, 53A17; Secondary 53A10, 70B10
1. Introduction
In an interesting paper [7], Pamfilos proved that a compact orientable surface in $\mathbb{R}^{3}$ is a sphere if and only if, for a sufficiently small ball rolling freely without skidding on the surface, the speed of its center is constant for all initial conditions; see also [8].
This result links geometry and dynamics. Its proof is based on the d’Alembert–Lagrange principle (with Lagrange multipliers), which the author uses to derive the equations of motion. It should be noted that the ball is allowed to spin while rolling.
Inspired by Pamfilos’ paper, here we obtain similar links between the geometry of a connected orientable surface $S$ and the kinematics of rolling motions, which can be proved using surface theory alone. The main difference is that, in absence of equations of motion, to define a “kinematic rolling” one must prescribe the entire locus of contact between the two surfaces, not just an initial condition.
To state our main result, fix $p\in S$, let $B_{r\neq 0}$ be the ball of radius $\lvert r\rvert$ centered at $p+rN_{p}$, where $N$ is the surface unit normal, and let $\gamma$ be a curve in $S$ such that $\gamma(0)=p$. Under a mild condition on $\gamma$, there exists a unique rolling (without skidding or spinning) of $B_{r}$ on $S$ such that $\gamma(t)$ is a point of contact with $B_{r}$; see Theorem 2.1. The speed of $B_{r}$ while rolling on $S$ is the length of the velocity of its center. We say that $B_{r}$ rolls isotropically from $p$ if its initial speed while rolling along $\gamma$ does not depend on the choice of direction $v\coloneqq\gamma^{\prime}(0)\in T_{p}S$.
Theorem 1.1.
If for some $r$ the initial speed of $B_{r}$ is the same for three nonparallel directions $v\in T_{p}S$, then $S$ is either umbilic or its mean curvature (with respect to $N$) equals $1/r$ at $p$. Conversely, if $S$ has mean curvature $h\neq 0$ and is umbilic (resp., nonumbilic) at $p$, then $B_{r}$ rolls isotropically from $p$ for all $r\neq 1/h$ (resp., only for $r=1/h$).
Remark 1.2.
If $p$ is umbilic and the mean curvature is $h$, then there is no rolling of $B_{1/h}$ on $S$.
We emphasize that while we use the term “ball” and the symbol $B_{r}$ for physical intuition and notational convenience, technically it is the sphere bounding $B_{r}$ that rolls on $S$; see section 2. Further note that in our setting intersections between $B_{r}$ and $S$ are allowed.
As direct consequences of the above theorem, we obtain the following mechanical characterizations of constant-mean-curvature (CMC) surfaces and spheres.
Corollary 1.3.
A connected orientable surface $S$ has nonzero constant mean curvature if and only if for some $r$ the ball $B_{r}$ rolls isotropically from every point of $S$.
Corollary 1.4.
Suppose that $S$ has bounded principal curvatures $\kappa_{1},\kappa_{2}$, and choose any $r\neq 2/(\kappa_{1}+\kappa_{2})$ such that $B_{r}$ rolls from every point of $S$. Then $S$ is a sphere if and only if $B_{r}$ rolls isotropically from every point of $S$.
Remark 1.5.
The conditions on $r$ in the last corollary are satisfied when $\lvert r\rvert<1/\lambda$ and $\lambda\coloneqq\sup_{p\in S,i=1,2}\lvert\kappa_{i}(p)\rvert$; or, replacing $\sup$ with $\inf$, when $\lvert r\rvert>1/\lambda$. Each of these inequalities guarantees that the normal curvature of any curve on (the boundary of) $B_{r}$ is never equal to that of $\gamma$, so that the rolling is well defined.
In the case of spheres, not only is the speed of $B_{r}$ constant with respect to the direction $v\in T_{p}S$, but also with respect to $p$. Applying Theorem 1.1, we can identify all surfaces with this property.
Corollary 1.6.
The only connected orientable surfaces where for some $r$ the ball $B_{r}$ rolls with constant speed along every curve are (subsets of) planes, circular cylinders, and spheres. In particular, $r=2R$ for a cylinder of radius $R$ oriented inward.
Theorem 1.1 (and Corollary 1.6) will be proved in section 3 on the basis of results of [6, 9], which we review in section 2. More explicitly, the proof hinges on the fact that, as $B_{r}$ rolls on $S$ along $\gamma$, the angular velocity vector is completely determined by the Darboux frames of $\gamma$ and its “anti-development” $\widetilde{\gamma}$ on $B_{r}$ (uniquely defined by the geodesic curvature of $\gamma$). This fact allows us to express the speed of $B_{r}$ solely in terms of $\lvert r\rvert$ and the Darboux curvatures of $\gamma$, and reduces the proof to a computation involving the principal curvatures of $S$.
It is worth noting that Theorem 1.1 can be generalized by allowing the rolling surface $\widetilde{S}$ to be different than a ball, as long as we interpret the speed of $\widetilde{S}$ as the length of its angular velocity vector, and we position $\widetilde{S}$ so that $v\in T_{p}S$ is a fixed principal direction $\widetilde{e}_{i}$ of $\widetilde{S}$. Then a minor modification of our proof yields that $\widetilde{S}$ rolls isotropically from $p$ if and only if either $p$ is umbilic or the mean curvature of $S$ equals $\widetilde{\kappa}_{i}$, where $\widetilde{\kappa}_{i}$ is the principal curvature corresponding to $\widetilde{e}_{i}$. This more general statement implies that a developable surface can roll isotropically on a minimal surface, while we already know that a ball cannot.
We conclude this introduction with some remarks on the literature. The problem of rolling two surfaces, or more generally two submanifolds, one on the other has a rich history. First considered by Chaplygin in 1897 [3], it was later studied by Nomizu [6], Bryant–Hsu [2, section 4], Sharpe [11, Appendix B], and Agrachev–Sachkov [1, Chapter 24], among others. Of particular interest nowadays is the intrinsic viewpoint, first presented in [5]. For a detailed historical perspective, discussing both classical and modern results, we refer the reader to the excellent survey [4].
2. Kinematic rolling
In this section we discuss some preliminaries. In the first part we follow rather closely [6]. Relevant textbook references for our discussion are [13, section 7.9] and [10, section 3.5].
Let $t\in[0,\ell]$. Consider a smooth one-parameter family $f_{t}$ of direct isometries $x\mapsto A_{t}x+b_{t}$ of $\mathbb{R}^{3}$ such that $f_{0}=\operatorname{id}$, and let $X_{t}$ be the time-dependent vector field $x\mapsto f_{t}^{\prime}(f_{t}^{-1}(x))$. We call $X_{t}$ the instantaneous motion at instant $t$. A computation shows that
$f_{t}^{\prime}(f_{t}^{-1}(x))=Q_{t}x+v_{t},$
where $Q_{t}\coloneqq A_{t}^{\prime}A_{T}^{T}$ and $v_{t}\coloneqq b^{\prime}_{t}-Q_{t}b_{t}$. In particular, the instantaneous motion is called
  1. (1)
    an instantaneous standstill if $Q_{t}=0$ and $b_{t}=0$;
  2. (2)
    an instantaneous translation if $Q_{t}=0$ and $b_{t}\neq 0$;
  3. (3)
    an instantaneous rotation if $Q_{t}\neq 0$ and there is a point $x_{0}$, called center, such that $f_{t}^{\prime}(f_{t}^{-1}(x_{0}))=0$.
Note that if $X_{t}$ is an instantaneous rotation, then there exists a unique vector $\omega_{t}$, the angular velocity vector, such that
$Q_{t}x=\omega_{t}\times x\quad\forall x\in\mathbb{R}^{3}.$
We are now ready to define rolling. Let $S$ and $\widetilde{S}$ be connected $\mathcal{C}^{2}$ surfaces tangent to each other at a point $p\in S$, and let $\gamma$ be a unit-speed curve on $S$ such that $\gamma(0)=p$. A motion of $\widetilde{S}$ on $S$ along $\gamma$ is a map $f_{t}$ with the property that $f_{t}(\widetilde{S})$ is tangent to $S$ at $\gamma(t)$. A motion of $\widetilde{S}$ is said to be rotational if $X_{t}$ is an instantaneous rotation with $\gamma(t)$ as center. A rotational motion is a rolling (without skidding or spinning) if the angular velocity vector is tangent to $S$ at $\gamma(t)$.
To construct a rolling of $\widetilde{S}$ on $S$ along $\gamma$, we proceed essentially as in [9]. Let $N$ be a unit vector field normal to $S$ along $\gamma$, and let $D_{t}$ be the matrix whose columns are the elements of the Darboux frame $(\gamma^{\prime}(t),N(t)\times\gamma^{\prime}(t),N(t))$ of $\gamma$. For any curve $\widetilde{\gamma}$ on $\widetilde{S}$ such that $\widetilde{\gamma}(0)=p$ and $\widetilde{\gamma}^{\prime}(0)=\gamma^{\prime}(0)$, set
$A_{t}$ $\coloneqq D_{t}\widetilde{D}_{t}^{T},$
$b_{t}$ $\coloneqq\gamma(t)-A_{t}\widetilde{\gamma}(t),$
where the matrix $\widetilde{D}_{t}$ is defined analogously, with $\widetilde{N}(0)=N(0)$. Then $f_{t}$ is a motion of $\widetilde{S}$ on $S$ along $\gamma$. Since
$f_{t}^{\prime}(f_{t}^{-1}(x))=Q_{t}(x-\gamma(t)),$
we observe that $f_{t}$ is rotational when $Q_{t}\neq 0$. In particular, we may compute that $Q_{t}=D_{t}^{\prime}D_{t}^{T}+D_{t}\widetilde{D}_{t}^{\prime T}\widetilde{D}_{t}D_{t}^{T}$. Next let
$\varLambda_{t}\coloneqq\begin{bmatrix}0&\kappa_{g}(t)&\kappa_{n}(t)\\ -\kappa_{g}(t)&0&\tau_{g}(t)\\ -\kappa_{n}(t)&-\tau_{g}(t)&0\end{bmatrix},$
where $\kappa_{g}\coloneqq\langle\gamma^{\prime\prime},N\times\gamma^{\prime}\rangle$, $\kappa_{n}\coloneqq\langle\gamma^{\prime\prime},N\rangle$, and $\tau_{g}\coloneqq-\langle N^{\prime},N\times\gamma^{\prime}\rangle$ are the geodesic curvature, the normal curvature, and the geodesic torsion of $\gamma$. Since $D_{t}^{\prime}=D_{t}\varLambda_{t}^{T}$ and $\varLambda_{t}$ is skew-symmetric, we obtain $Q_{t}=D_{t}(\widetilde{\varLambda}_{t}-\varLambda_{t})D_{t}^{T}$. Hence
$\omega_{t}=(\tau_{g}(t)-\widetilde{\tau}_{g}(t))\gamma^{\prime}(t)-(\kappa_{n}(t)-\widetilde{\kappa}_{n}(t))N(t)\times\gamma^{\prime}(t)+(\kappa_{g}(t)-\widetilde{\kappa}_{g}(t))N(t).$
The following theorem summarizes our findings.
Theorem 2.1 ([6, 9]).
Let $\widetilde{\gamma}^{\ast}$ be the unique curve $\widetilde{\gamma}$ whose geodesic curvature coincides with $\kappa_{g}$. There exists a (unique) rolling of $\widetilde{S}$ on $S$ along $\gamma$ if and only if $\kappa_{n}$ and $\tau_{g}$ are never simultaneously equal to $\widetilde{\kappa}_{n}^{\ast}$ and $\widetilde{\tau}_{g}^{\ast}$. Moreover, if $\widetilde{S}$ rolls on $S$ along $\gamma$, then the angular velocity has components $(\tau_{g}-\widetilde{\tau}_{g}^{\ast},\widetilde{\kappa}_{n}^{\ast}-\kappa_{n},0)$ in the Darboux frame of $\gamma$.
3. Proof of the main result
Here we prove Theorem 1.1 and Corollary 1.6.
Let $h$ be the mean curvature of $S$ at $p$. We begin by showing that when $S$ is nonumbilic, the rolling of $B_{1/h}$ is well defined in a neighborhood of $p$.
Lemma 3.1.
If $p$ is not an umbilic point, then for any $v\in T_{p}S$ and curve $\gamma\colon[0,\ell]\to S$ such that $\gamma(0)=p$ and $\gamma^{\prime}(0)=v$, there exists an $\varepsilon>0$ and a unique rolling of $B_{1/h}$ on $S$ along $\gamma\rvert_{[0,\varepsilon]}$.
Proof.
It is clear that any curve on $B_{1/h}$ (oriented inward) has normal curvature $h$ and zero geodesic torsion. Hence, by Theorem 2.1, we need to show that if $\kappa_{n}(0)=h$ for some $v\in T_{p}S$, then $\tau_{g}(0)\neq 0$. This will imply that $\tau_{g}(t)\neq 0$ in a neighborhood of 00, as desired.
Let $\kappa_{1}\geq\kappa_{2}$ be the principal curvatures of $S$ at $p$, and let $e_{1},e_{2}$ be a positively oriented basis of principal unit vectors. Given $v\in T_{p}S$, denote by $\theta$ the angle between $v$ and $e_{1}$. Then, by Euler’s theorem,
(1) $\kappa_{n}(\theta)$ $=\kappa_{1}\cos^{2}\theta+\kappa_{2}\sin^{2}\theta,$
(2) $\tau_{g}(\theta)$ $=(\kappa_{2}-\kappa_{1})\sin\theta\cos\theta.$
Suppose that $\kappa_{n}(\theta)=1/2(\kappa_{1}+\kappa_{2})$. Then $\theta=\pm\pi/4$ or $\theta=\pm 3\pi/4$, which implies $\tau_{g}(\theta)=\pm(\kappa_{1}-\kappa_{2})/2$. Hence $\tau_{g}(\theta)$ can only vanish when $p$ is an umbilic point, and the lemma is proved. ∎
Proof of Theorem 1.1.
By Theorem 2.1, we know that as $B_{r}$ rolls on $S$ along $\gamma$, the angular velocity at instant $t$ is given by
$\omega_{t}=\tau_{g}(t)\gamma^{\prime}(t)+(1/r-\kappa_{n}(t))N(t)\times\gamma^{\prime}(t).$
The velocity of the center of $B_{r}$ is thus
$\omega_{t}\times rN(t)=(1-r\kappa_{n}(t))\gamma^{\prime}(t)-r\tau_{g}(t)N(t)\times\gamma^{\prime}(t),$
which has length squared
(3) $(1-r\kappa_{n}(t))^{2}+r^{2}\tau_{g}^{2}(t).$
In particular, for $t=0$, we may substitute equations (1) and (2) into (3). This gives
$\left(1-r\left(\kappa_{1}\cos^{2}\theta+\kappa_{2}\sin^{2}\theta\right)\right)^{2}+r^{2}\left((\kappa_{2}-\kappa_{1})\sin\theta\cos\theta\right)^{2}\eqqcolon\ell(\theta).$
Expanding the squares and using the Pythagorean identity, we obtain
$\ell(\theta)=1+r^{2}\left(\kappa_{2}^{2}-\kappa_{2}^{2}\cos^{2}\theta+\kappa_{1}^{2}\cos^{2}\theta\right)-2r\left(\kappa_{1}\cos^{2}\theta+\kappa_{2}-\kappa_{2}\cos^{2}\theta\right),$
which we rewrite as
(4) $\ell(\theta)=1+r^{2}\kappa_{2}^{2}-2r\kappa_{2}+r(\kappa_{2}-\kappa_{1})(2-r(\kappa_{1}+\kappa_{2}))\cos^{2}\theta.$
Since $\cos^{2}$ is $\pi$-periodic, this equation implies that if the initial speed is the same for more than four values of $\theta$, then it does not depend on $\theta$. Moreover, the latter condition occurs if and only if
$(\kappa_{2}-\kappa_{1})(2-r(\kappa_{1}+\kappa_{2}))=0,$
i.e., when either $p$ is umbilic or $r=1/h$, which is the desired conclusion. ∎
Proof of Corollary 1.6.
By equation (3), it is clear that planes and spheres are examples. We may therefore assume that $S$ is free of umbilics in a neighborhood $U$ of $p$. Then both the principal curvatures $\kappa_{1}>\kappa_{2}$ and principal directions $e_{1},e_{2}$ depend continuously on $p\in U$; see [12].
Let $r=2/(\kappa_{1}+\kappa_{2})$. Then (4) reads
$\ell(\theta)=1+\frac{4\kappa_{2}^{2}}{(\kappa_{1}+\kappa_{2})^{2}}-\frac{4\kappa_{2}}{\kappa_{1}+\kappa_{2}}=\frac{(\kappa_{1}-\kappa_{2})^{2}}{(\kappa_{1}+\kappa_{2})^{2}}.$
Since $\kappa_{1}+\kappa_{2}$ is constant by assumption, it follows that the speed of $B_{r}$ can only be constant if both principal curvatures are constant. But, as it is well known, if a surface has constant principal curvatures, then it is (part of) a plane, a sphere, or a circular cylinder. Hence $U$ must be part of a circular cylinder, as desired. ∎
Acknowledgments
The author thanks Steen Markvorsen and Vicente Palmer for interesting discussions.

### v3

@@
 Abstract.
-The speed of a ball rolling without skidding or spinning on a surface $S$ is the length of the velocity of its center, or, equivalently, the length of its angular velocity vector.
-We show that if the speed only depends on $p\in S$, then $S$ has nonzero constant mean curvature; and, conversely, that if the mean curvature of $S$ is constant and equal to $H\neq 0$, then either $S$ is a sphere or the ball of radius $1/H$ rolls on $S$ with direction-independent speed.
+The speed of a ball rolling without skidding or spinning on a surface $S$ is the length of the velocity of its center.
+We show that if the speed depends only on $p\in S$, then $S$ has constant mean curvature; and, conversely, that if the mean curvature of $S$ is constant and equal to $H\neq 0$, then either $S$ is a sphere or the ball of radius $1/H$ rolls on $S$ with direction-independent speed.
 It follows that the only surfaces where the speed is constant are subsets of planes, circular cylinders, and spheres.
@@
 1. Introduction
-In an interesting paper [7], Pamfilos proved that a compact orientable surface in $\mathbb{R}^{3}$ is a sphere if and only if, for a sufficiently small ball rolling freely without skidding on the surface, the speed of its center is constant for all initial conditions; see also [8].
+In an interesting paper [8], Pamfilos proved that a compact orientable surface in $\mathbb{R}^{3}$ is a sphere if and only if, given a sufficiently small ball rolling freely without skidding on the surface, the ball’s center moves with constant speed for any initial condition; see also [9].
 This result links geometry and dynamics.
-Its proof is based on the d’Alembert–Lagrange principle (with Lagrange multipliers), which the author uses to derive the equations of motion.
+Its proof is based on the d’Alembert–Lagrange principle, which the author uses to derive the equations of motion.
 It should be noted that the ball is allowed to spin while rolling.
-Inspired by Pamfilos’ paper, here we obtain similar links between the geometry of a connected orientable surface $S$ and the kinematics of rolling motions, which can be proved using surface theory alone.
-The main difference is that, in absence of equations of motion, to define a “kinematic rolling” one must prescribe the entire locus of contact between the two surfaces, not just an initial condition.
+Inspired by Pamfilos’ paper, here we obtain similar links between the geometry of a connected orientable surface $S$ and the kinematics of rolling motions, using surface theory alone.
+Following Nomizu’s viewpoint [7], a rolling (without skidding or spinning) of one surface on another is described extrinsically by a one-parameter family of rigid motions that keeps the moving surface tangent to $S$ along a prescribed contact curve $\gamma$.
+The no-skid condition means that the instantaneous relative velocity vanishes at the point of contact, while the no-spin condition means that the instantaneous rotation does not include any component about the common normal.
+Thus, in the absence of equations of motion, to define a kinematic rolling one must prescribe the entire locus of contact $\gamma$ between the two surfaces, rather than merely an initial condition.
 To state our main result, fix $p\in S$, let $B_{r\neq 0}$ be the ball of radius $\lvert r\rvert$ centered at $p+rN_{p}$, where $N$ is the surface unit normal, and let $\gamma$ be a curve in $S$ such that $\gamma(0)=p$.
@@
 Theorem 1.1.
-If for some $r$ the initial speed of $B_{r}$ is the same for three nonparallel directions $v\in T_{p}S$, then $S$ is either umbilic or its mean curvature (with respect to $N$) equals $1/r$ at $p$.
+If for some $r$ the initial speed of $B_{r}$ is the same for three pairwise nonparallel directions $v\in T_{p}S$, then $S$ is either umbilic or its mean curvature (with respect to $N$) equals $1/r$ at $p$.
 Conversely, if $S$ has mean curvature $h\neq 0$ and is umbilic (resp., nonumbilic) at $p$, then $B_{r}$ rolls isotropically from $p$ for all $r\neq 1/h$ (resp., only for $r=1/h$).
 Remark 1.2.
-If $p$ is umbilic and the mean curvature is $h$, then there is no rolling of $B_{1/h}$ on $S$.
-We emphasize that while we use the term “ball” and the symbol $B_{r}$ for physical intuition and notational convenience, technically it is the sphere bounding $B_{r}$ that rolls on $S$; see section 2.
-Further note that in our setting intersections between $B_{r}$ and $S$ are allowed.
-As direct consequences of the above theorem, we obtain the following mechanical characterizations of constant-mean-curvature (CMC) surfaces and spheres.
+If $p$ is umbilic and the mean curvature is $h\neq 0$, then there is no rolling of $B_{1/h}$ on $S$.
+Indeed, in that case the normal curvature and geodesic torsion of $\gamma$ at $\gamma(0)$ coincide with the corresponding quantities on the ball, which violates the existence condition in Theorem 2.1.
+We emphasize that, although we use the term “ball” and the symbol $B_{r}$, it is technically the sphere bounding $B_{r}$ that rolls on $S$; see section 2.
+We further note that, in our setting, intersections between $B_{r}$ and $S$ are allowed.
+As a consequence of the above theorem, we obtain the following mechanical characterizations of nonminimal constant-mean-curvature (CMC) surfaces and spheres.
 Corollary 1.3.
-A connected orientable surface $S$ has nonzero constant mean curvature if and only if for some $r$ the ball $B_{r}$ rolls isotropically from every point of $S$.
+Suppose that $S$ is not totally umbilic. Then $S$ has nonzero constant mean curvature if and only if for some $r$ the ball $B_{r}$ rolls isotropically from every nonumbilic point of $S$.
+Proof.
+Let $U$ denote the set of umbilic points of $S$, and suppose that $B_{r}$ rolls isotropically from every point of $S\setminus U$.
+Then the mean curvature $H$ equals $1/r$ on $S\setminus U$.
+By continuity, $H=1/r$ on the closure $\overline{S\setminus U}$.
+Now let $W$ be any connected component of $\operatorname{int}(U)$.
+Since every point of $W$ is umbilic, $W$ lies in a sphere or a plane, so $H$ is constant on $W$.
+Because $\partial W\subset\overline{S\setminus U}$, continuity implies $H=1/r$ on $W$ as well, completing the proof.
+∎
+Remark 1.4.
+On a connected CMC surface, umbilic points are isolated unless the surface is totally umbilic [14, p. 239].
 Corollary 1.5.
-Suppose that $S$ has bounded principal curvatures $\kappa_{1},\kappa_{2}$, and choose any $r\neq 2/(\kappa_{1}+\kappa_{2})$ such that $B_{r}$ rolls from every point of $S$.
+Suppose that $S$ is compact, and choose $r$ such that $1/\lvert r\rvert>\max_{p\in S,i=1,2}\lvert\kappa_{i}(p)\rvert$, where $\kappa_{1},\kappa_{2}$ are the principal curvatures.
 Then $S$ is a sphere if and only if $B_{r}$ rolls isotropically from every point of $S$.
 Remark 1.6.
-The conditions on $r$ in the last corollary are satisfied when $\lvert r\rvert<1/\lambda$ and $\lambda\coloneqq\sup_{p\in S,i=1,2}\lvert\kappa_{i}(p)\rvert$; or, replacing $\sup$ with $\inf$, when $\lvert r\rvert>1/\lambda$.
-Each of these inequalities guarantees that the normal curvature of any curve on (the boundary of) $B_{r}$ is never equal to that of $\gamma$, so that the rolling is well defined.
+The condition on $r$ in the last corollary ensures that the normal curvature of any curve on (the boundary of) $B_{r}$ never coincides with that of $\gamma$, and that $(\kappa_{1}(p)+\kappa_{2}(p))r\neq 2$ for all $p\in S$.
 In the case of spheres, not only is the speed of $B_{r}$ constant with respect to the direction $v\in T_{p}S$, but also with respect to $p$.
@@
 Corollary 1.7.
-The only connected orientable surfaces where for some $r$ the ball $B_{r}$ rolls with constant speed along every curve are (subsets of) planes, circular cylinders, and spheres.
+The only connected orientable surfaces where for some $r$ the speed of the ball $B_{r}$ is constant are (subsets of) planes, circular cylinders, and spheres.
 In particular, $r=2R$ for a cylinder of radius $R$ oriented inward.
 Theorem 1.1 (and Corollary 1.7) will be proved in section 3 on the basis of results of [7, 10], which we review in section 2.
-More explicitly, the proof hinges on the fact that, as $B_{r}$ rolls on $S$ along $\gamma$, the angular velocity vector is completely determined by the Darboux frames of $\gamma$ and its “anti-development” $\widetilde{\gamma}$ on $B_{r}$ (uniquely defined by the geodesic curvature of $\gamma$).
+More explicitly, the proof relies on the fact that, as $B_{r}$ rolls on $S$ along $\gamma$, the angular velocity vector is completely determined by the Darboux frames of $\gamma$ and its “anti-development” $\widetilde{\gamma}$ on $B_{r}$ (the locus of contact on $B_{r}$, uniquely defined by the geodesic curvature of $\gamma$).
 This fact allows us to express the speed of $B_{r}$ solely in terms of $\lvert r\rvert$ and the Darboux curvatures of $\gamma$, and reduces the proof to a computation involving the principal curvatures of $S$.
-It is worth noting that Theorem 1.1 can be generalized by allowing the rolling surface $\widetilde{S}$ to be different than a ball, as long as we interpret the speed of $\widetilde{S}$ as the length of its angular velocity vector, and we position $\widetilde{S}$ so that $v\in T_{p}S$ is a fixed principal direction $\widetilde{e}_{i}$ of $\widetilde{S}$.
-Then a minor modification of our proof yields that $\widetilde{S}$ rolls isotropically from $p$ if and only if either $p$ is umbilic or the mean curvature of $S$ equals $\widetilde{\kappa}_{i}$, where $\widetilde{\kappa}_{i}$ is the principal curvature corresponding to $\widetilde{e}_{i}$.
-This more general statement implies that a developable surface can roll isotropically on a minimal surface, while we already know that a ball cannot.
+It is worth noting that Theorem 1.1 can be generalized by allowing the rolling surface $\widetilde{S}$ to be different from a ball, as long as we interpret the speed of $\widetilde{S}$ as the length of its angular velocity vector, and we position $\widetilde{S}$ so that $v\in T_{p}S$ is a fixed principal direction $\widetilde{e}_{i}$ of $\widetilde{S}$.
+Then a minor modification of the proof yields that $\widetilde{S}$ rolls isotropically from $p$ if and only if either $p$ is umbilic or the mean curvature of $S$ equals $\widetilde{\kappa}_{i}$, where $\widetilde{\kappa}_{i}$ is the principal curvature corresponding to $\widetilde{e}_{i}$.
+This more general statement implies that a developable surface can roll isotropically on a nonplanar minimal surface, while we already know that a ball cannot.
 We conclude this introduction with some remarks on the literature.
 The problem of rolling two surfaces, or more generally two submanifolds, one on the other has a rich history.
-First considered by Chaplygin in 1897 [3], it was later studied by Nomizu [6], Bryant–Hsu [2, section 4], Sharpe [11, Appendix B], and Agrachev–Sachkov [1, Chapter 24], among others.
+First considered by Chaplygin in 1897 [4], it was later studied by Nomizu [7], Bryant–Hsu [3, section 4], Sharpe [12, Appendix B], and Agrachev–Sachkov [1, Chapter 24], among others.
 Of particular interest nowadays is the intrinsic viewpoint, first presented in [6].
@@
 2. Kinematic rolling
-In this section we discuss some preliminaries. In the first part we follow rather closely [6]. Relevant textbook references for our discussion are [13, section 7.9] and [10, section 3.5].
+In this section we discuss some preliminaries. In the first part we follow rather closely [7]. Relevant textbook references are [14, section 7.9] and [11, section 3.5].
 Let $t\in[0,\ell]$.
@@
   1. (1)
-    an instantaneous standstill if $Q_{t}=0$ and $b_{t}=0$;
+    an instantaneous standstill if $Q_{t}=0$ and $v_{t}=0$;
   2. (2)
-    an instantaneous translation if $Q_{t}=0$ and $b_{t}\neq 0$;
+    an instantaneous translation if $Q_{t}=0$ and $v_{t}\neq 0$;
   3. (3)
@@
 $Q_{t}x=\omega_{t}\times x\quad\forall x\in\mathbb{R}^{3}.$
-We are now ready to define rolling.
-Let $S$ and $\widetilde{S}$ be connected $\mathcal{C}^{2}$ surfaces tangent to each other at a point $p\in S$, and let $\gamma$ be a unit-speed curve on $S$ such that $\gamma(0)=p$.
-A motion of $\widetilde{S}$ on $S$ along $\gamma$ is a map $f_{t}$ with the property that $f_{t}(\widetilde{S})$ is tangent to $S$ at $\gamma(t)$.
-A motion of $\widetilde{S}$ is said to be rotational if $X_{t}$ is an instantaneous rotation with $\gamma(t)$ as center.
-A rotational motion is a rolling (without skidding or spinning) if the angular velocity vector is tangent to $S$ at $\gamma(t)$.
+[…]
 To construct a rolling of $\widetilde{S}$ on $S$ along $\gamma$, we proceed essentially as in [10].
 Let $N$ be a unit vector field normal to $S$ along $\gamma$, and let $D_{t}$ be the matrix whose columns are the elements of the Darboux frame $(\gamma^{\prime}(t),N(t)\times\gamma^{\prime}(t),N(t))$ of $\gamma$.
-For any curve $\widetilde{\gamma}$ on $\widetilde{S}$ such that $\widetilde{\gamma}(0)=p$ and $\widetilde{\gamma}^{\prime}(0)=\gamma^{\prime}(0)$, set
+For any curve $\widetilde{\gamma}\colon[0,\ell]\to\widetilde{S}$ such that $\widetilde{\gamma}(0)=p$ and $\widetilde{\gamma}^{\prime}(0)=\gamma^{\prime}(0)$, let $\widetilde{D}_{t}$ be defined analogously by the normal $\widetilde{N}$ that coincides with $N$ at $p$, so that $D_{0}=\widetilde{D}_{0}$.
+Then, setting
 $A_{t}$ $\coloneqq D_{t}\widetilde{D}_{t}^{T},$
 $b_{t}$ $\coloneqq\gamma(t)-A_{t}\widetilde{\gamma}(t),$
-where the matrix $\widetilde{D}_{t}$ is defined analogously, with $\widetilde{N}(0)=N(0)$. Then $f_{t}$ is a motion of $\widetilde{S}$ on $S$ along $\gamma$. Since
+the family $f_{t}$ becomes a motion of $\widetilde{S}$ on $S$ along $\gamma$ [10, Prop. 2.2]. Since
 $f_{t}^{\prime}(f_{t}^{-1}(x))=Q_{t}(x-\gamma(t)),$
@@
 $\omega_{t}=(\tau_{g}(t)-\widetilde{\tau}_{g}(t))\gamma^{\prime}(t)-(\kappa_{n}(t)-\widetilde{\kappa}_{n}(t))N(t)\times\gamma^{\prime}(t)+(\kappa_{g}(t)-\widetilde{\kappa}_{g}(t))N(t).$
-The following theorem summarizes our findings.
+The following theorem concludes our discussion.
 Theorem 2.1 ([7, 10]).
-Let $\widetilde{\gamma}^{\ast}$ be the unique curve $\widetilde{\gamma}$ whose geodesic curvature coincides with $\kappa_{g}$.
-There exists a (unique) rolling of $\widetilde{S}$ on $S$ along $\gamma$ if and only if $\kappa_{n}$ and $\tau_{g}$ are never simultaneously equal to $\widetilde{\kappa}_{n}^{\ast}$ and $\widetilde{\tau}_{g}^{\ast}$.
+Suppose that $\widetilde{S}$ is complete, and let $\widetilde{\gamma}^{\ast}$ be the unique curve $\widetilde{\gamma}$ whose geodesic curvature coincides with $\kappa_{g}$.
+Then there exists a (unique) rolling of $\widetilde{S}$ on $S$ along $\gamma$ if and only if $\kappa_{n}$ and $\tau_{g}$ are never simultaneously equal to $\widetilde{\kappa}_{n}^{\ast}$ and $\widetilde{\tau}_{g}^{\ast}$.
 Moreover, if $\widetilde{S}$ rolls on $S$ along $\gamma$, then the angular velocity has components $(\tau_{g}-\widetilde{\tau}_{g}^{\ast},\widetilde{\kappa}_{n}^{\ast}-\kappa_{n},0)$ in the Darboux frame of $\gamma$.
+Remark 2.2.
+In the introduction we referred to the curve $\widetilde{\gamma}^{\ast}$ as the anti-development of $\gamma$.
+When $\widetilde{S}$ is complete, standard ODE theory guarantees that $\widetilde{\gamma}^{\ast}$ exists and is unique; see [2, Thm.
+2.1] for a proof.
+We emphasize that, although $B_{r}$ is complete, for our purposes it suffices to roll along an arbitrarily short subarc of $\gamma$.
 3. Proof of the main result
 Here we prove Theorem 1.1 and Corollary 1.7.
-Let $h$ be the mean curvature of $S$ at $p$. We begin by showing that when $S$ is nonumbilic, the rolling of $B_{1/h}$ is well defined in a neighborhood of $p$.
+Let $h$ be the mean curvature of $S$ at $p$. We begin by showing that when $h\neq 0$ and $S$ is nonumbilic, the rolling of $B_{1/h}$ is well defined in a neighborhood of $p$.
 Lemma 3.1.
-If $p$ is not an umbilic point, then for any $v\in T_{p}S$ and curve $\gamma\colon[0,\ell]\to S$ such that $\gamma(0)=p$ and $\gamma^{\prime}(0)=v$, there exists an $\varepsilon>0$ and a unique rolling of $B_{1/h}$ on $S$ along $\gamma\rvert_{[0,\varepsilon]}$.
+If $h\neq 0$ and $p$ is not an umbilic point, then for any $v\in T_{p}S$ and curve $\gamma\colon[0,\ell]\to S$ such that $\gamma(0)=p$ and $\gamma^{\prime}(0)=v$, there exists an $\varepsilon>0$ and a unique rolling of $B_{1/h}$ on $S$ along $\gamma\rvert_{[0,\varepsilon]}$.
 Proof.
-It is clear that any curve on $B_{1/h}$ (oriented inward) has normal curvature $h$ and zero geodesic torsion.
+It is clear that any curve on $B_{1/h}$ (oriented by $N_{p}$) has normal curvature $h$ and zero geodesic torsion.
 Hence, by Theorem 2.1, we need to show that if $\kappa_{n}(0)=h$ for some $v\in T_{p}S$, then $\tau_{g}(0)\neq 0$.
-This will imply that $\tau_{g}(t)\neq 0$ in a neighborhood of 00, as desired.
+This will imply that $\tau_{g}(t)\neq 0$ in a neighborhood of 0, as desired.
 Let $\kappa_{1}\geq\kappa_{2}$ be the principal curvatures of $S$ at $p$, and let $e_{1},e_{2}$ be a positively oriented basis of principal unit vectors.
@@
 (4) $\ell(\theta)=1+r^{2}\kappa_{2}^{2}-2r\kappa_{2}+r(\kappa_{2}-\kappa_{1})(2-r(\kappa_{1}+\kappa_{2}))\cos^{2}\theta.$
-Since $\cos^{2}$ is $\pi$-periodic, this equation implies that if the initial speed is the same for more than four values of $\theta$, then it does not depend on $\theta$.
-Moreover, the latter condition occurs if and only if
+Since $\cos^{2}\theta=k$ has at most four solutions on $0,2\pi]$ and $\ell(\theta)=\ell(\theta+\pi)$, equation (4) shows that if the initial speed is the same for at least three pairwise nonparallel directions, then it must be independent of $\theta$.
+Moreover, this occurs if and only if
 $(\kappa_{2}-\kappa_{1})(2-r(\kappa_{1}+\kappa_{2}))=0,$
-i.e., when either $p$ is umbilic or $r=1/h$, which is the desired conclusion. ∎
+that is, either $p$ is umbilic or $r=1/h$, which is the desired conclusion. ∎
 Proof of Corollary 1.7.
 By equation (3), it is clear that planes and spheres are examples.
-We may therefore assume that $S$ is free of umbilics in a neighborhood $U$ of $p$.
-Then both the principal curvatures $\kappa_{1}>\kappa_{2}$ and principal directions $e_{1},e_{2}$ depend continuously on $p\in U$; see [12].
-Let $r=2/(\kappa_{1}+\kappa_{2})$. Then (4) reads
+We may therefore assume that $S$ is free of umbilics in a neighborhood $V$ of $p$.
+Then the principal curvatures $\kappa_{1}>\kappa_{2}$ depend continuously on $p\in V$, and one can choose a continuous orthonormal principal frame $(e_{1},e_{2})$ on $V$; see [13].
+Suppose that the speed of $B_{r}$ is constant on $V$.
+By Corollary 1.3 and Theorem 1.1, it follows that $V$ has nonzero constant mean curvature and $r=2/(\kappa_{1}+\kappa_{2})$.
+Substituting into (4), we obtain
 $\ell(\theta)=1+\frac{4\kappa_{2}^{2}}{(\kappa_{1}+\kappa_{2})^{2}}-\frac{4\kappa_{2}}{\kappa_{1}+\kappa_{2}}=\frac{(\kappa_{1}-\kappa_{2})^{2}}{(\kappa_{1}+\kappa_{2})^{2}}.$
-Since $\kappa_{1}+\kappa_{2}$ is constant by assumption, it follows that the speed of $B_{r}$ can only be constant if both principal curvatures are constant.
+Since $\kappa_{1}+\kappa_{2}$ is constant, it follows that the speed of $B_{r}$ can only be constant if both principal curvatures are constant.
 But, as it is well known, if a surface has constant principal curvatures, then it is (part of) a plane, a sphere, or a circular cylinder.
-Hence $U$ must be part of a circular cylinder, as desired.
+Hence $V$ must be part of a circular cylinder, as desired.
 ∎